OpenStudy (anonymous):

P (one die was 2 | the sum is 6)

jimthompson5910 (jim_thompson5910):

Given this list of numbers 1, 2, 3, 4, 5, 6 how many ways are there to add any two numbers (in that list above) to 6?

OpenStudy (anonymous):

If one die was already 2, there would only be one possible answer. 6-2=?

OpenStudy (anonymous):

Right?

jimthompson5910 (jim_thompson5910):

List all the ways to add to 6 1+5 = 6 2+4 = 6 3+3 = 6 4+2 = 6 5+1 = 6 how many of these ways have a 2 in them?

OpenStudy (anonymous):

Yes but.... 2+4 is the same as 4+2 so there is only one possible answer.